# Supervised Manifold Learning for Functional Data

### v1

@@ Abstract
Classification is a core topic in functional data analysis. A large number of functional classifiers have been proposed in the literature, most of which are based on functional principal component analysis or functional regression. In contrast, we investigate this topic from the perspective of manifold learning. It is assumed that functional data lie on an unknown low-dimensional manifold, and we expect that better classifiers can be built upon the manifold structure. To this end, we propose a novel proximity measure that takes the label information into account to learn the low-dimensional representations, also known as the supervised manifold learning outcomes. When the outcomes are coupled with multivariate classifiers, the procedure induces a family of new functional classifiers. In theory, we show that our functional classifier induced by the $k$-NN classifier is asymptotically optimal. In practice, we show that our method, coupled with several classical multivariate classifiers, achieves outstanding classification performance compared to existing functional classifiers in both synthetic and real data examples.
@@ 1 Introduction
Functional data analysis has become increasingly prevalent across diverse scientific fields, including medicine, neuroscience, astronomy, chemometrics, etc. As a result, there is a growing demand for sophisticated models and algorithms designed for functional data analysis. A core topic in functional data analysis is classification, i.e., train a functional classifier based on observed data to predict the label for future functional observation. There exists a fairly rich literature on this topic. Since functional data are infinite-dimensional, a large number of existing functional classifiers are built on the functional principal component analysis (FPCA), whose theoretical basis is due to the Karhunen–Loève expansion: A second order random process $X$ can be decomposed as
@@
where $\mu(t)=E\{X(t)\}$, $\zeta_{j}$ are the PC scores, $\phi_{j}$ are the eigenfunctions of the covariance operator, also referred to as the PC basis. A finite number of the PC scores serves as a finite-dimensional approximation of the random process. Delaigle and Hall, (2013) and Galeano et al., (2015) adapted the linear and quadratic discriminant analysis to the PC scores to develop functional linear and quadratic discriminant classifiers. Delaigle and Hall, (2012) proposed to project functional data onto a one-dimensional space spanned by a linear combination of the PC basis, and showed that asymptotic perfect classification can be achieved under certain conditions. Wu and Liu, (2013) performed the support vector machines on the PC scores. Dai et al., (2017) proposed to estimate the densities of the PC scores nonparametrically and to utilize the Bayes classifier. Wang et al., 2024b constructed a deep neural network on the PC scores as a functional classifier, and Wang et al., (2023) generalized it to the setting of multidimensional functional data. In addition to FPCA-based methods, other functional classifiers include regression-based methods (Ferraty and Vieu,, 2003; Leng and Müller,, 2006; Preda et al.,, 2007; Araki et al.,, 2009), distance-based methods (Alonso et al.,, 2012; Kraus and Stefanucci,, 2019), nearest neighbors (Biau et al.,, 2005, 2010); see Wang et al., 2024a for a recent review.
@@
In this article, we propose a novel manifold learning procedure for functional data by taking the label information into account, which induces a family of new functional classifiers. Instead of the FPCA or regression, we start from the assumption that functional data lie on an unknown low-dimensional manifold. This assumption for high-dimensional data forms the basis for manifold learning (Meilă and Zhang,, 2024), and has also been invoked under the functional data setting (Chen and Müller,, 2012; Lin and Yao,, 2021; Tan et al.,, 2024). A motivating example of the manifold assumption is that functional data often exhibit phase variation, which inflates the variance of data and distorts the PCs (Marron et al.,, 2015). To investigate the non-linearity of phase variation, manifold techniques have been developed (Kneip and Gasser,, 1992; Srivastava et al.,, 2011).
@@
Based on the manifold assumption, we propose a novel proximity measure that includes the label information to quantify distance between functional data. Not only does the proximity measure penalize for different classes, it also tries to maintain the underlying manifold structure in terms of the geodesic distance. The latter property is crucial to classification, because the manifold learning step being continuous is necessary for an interpolator to accurately predict the low-dimensional representation of a future observation; see Section 2.2 for more discussion. Once the proximity measure is obtained for all pair-wise individuals, we perform the multidimensional scaling (MDS, Cox and Cox,, 2008) to obtain the low-dimensional manifold learning outcomes. Such outcomes can also be used for data visualization if the target dimension is two or three. Finally, a smooth interpolator accompanied with any reasonable multivariate classifier yields a functional classifier. In theory, we provide the with-in-sample and out-of-sample convergence rates of the estimated coordinate map. As for classification, we also show that our functional classifier coupled with the vanilla $k$-NN classifier is asymptotically optimal. Note that the intrinsic dimension being low is opposite to the number of PCs being high when the manifold is flat. The latter is a core assumption to show nice theoretical properties, say, the asymptotic perfect classification, of several functional classifiers (Delaigle and Hall,, 2012; Dai et al.,, 2017; Berrendero et al.,, 2018). Therefore, our investigation of the low intrinsic dimension serves as an interesting complement to the existing literature.
@@
Manifold learning has been an active research field since the seminal works by Tenenbaum et al., (2000) and Roweis and Saul, (2000), which has gained tremendous success in bioinformatics (Moon et al.,, 2018), image processing (Pless and Souvenir,, 2009), etc. In functional data analysis, Chen and Müller, (2012) extended the Isomap (Tenenbaum et al.,, 2000) to functional data and developed the functional manifold component; Tan et al., (2024) extended the parallel transport unfolding (Budninskiy et al.,, 2019) to functional data and conducted a clustering analysis. However, their works belong to unsupervised manifold learning, i.e., no label variable is involved, and thus their methods are not suitable for classification. In contrast, our manifold learning procedure is designed to take the label information into account, also known as supervised manifold learning. There is also a vast literature on supervised manifold learning for high-dimensional data; see Chao et al., (2019) for a survey. Our method may be categorized into Isomap-based methods, but it differs from existing supervised Isomap methods (Vlachos et al.,, 2002; Geng et al.,, 2005; Zhang et al.,, 2018) mainly from two aspects: Our proximity measure is new, and we are dealing with functional data instead of high-dimensional data.
@@
The rest of this article is organized as follows. In Section 2.1, we first introduce the model and present two examples to show the difficulty of the problem and the motivation of our method. We then introduce our main method in Section 2.2. In Section 3, we discuss a detailed estimation procedure to implement the proposed method, including the selection of tuning parameters. We provide theoretical properties of our estimators in Section 4, and conduct simulation studies and real data applications in Sections 5 and 6, respectively. We conclude in Section 7. The Supplementary Material contains technical proofs and further implementation details.
@@ 2.1 Model and Synthetic Examples
It is reasonable to assume that the functional variables from different classes are concentrated in different regions of the manifold, i.e., the modes of the distributions $X|Y=y$ are easily separated in $\mathcal{M}$. However, the optimal boundaries in separating classes may be highly nonlinear in the original space so that standard linear classification methods may fail. Yet, the optimal boundaries may become simple if the data are properly embedded in a low-dimensional space. Example 1 illustrates this phenomenon in $\mathbb{R}^{3}$, and Example 2 is a more concrete example showing the classification difficulty under the functional manifold setting.
@@
Figure 2: The panels from left to right show a simulated sample of curves as in Example 2, the first two principal component scores (PC1 and PC2) using the FPCA, the unsupervised manifold learning outcomes $\widetilde{Z}=(\widetilde{Z}_{1},\widetilde{Z}_{2})$ using the FIsomap and the supervised manifold learning outcomes $Z=(Z_{1},Z_{2})$ using our proposed method FSML, respectively. The class $Y=1$ is denoted by solid line (--) and star (*), while $Y=2$ is denoted by dashed line (----) and circle ($\circ$).
@@
[…]
For the label variable $Y$, we set $Y=1$ if $w\in[0,\pi]$ and $Y=2$ otherwise. The first panel of Figure 2 shows a simulated sample of 200 curves (half in each class) with $g=\sin(2\pi t)$ and $k_{0}=50$. We apply the FPCA to the sample, whose first two principal component scores are shown in the second panel of Figure 2. The manifold learning outcomes using the FIsomap (Chen and Müller,, 2012) and our proposed functional supervised manifold learning (FSML) are shown in the third and fourth panels of Figure 2, respectively. We see that the PC scores and the FIsomap outcomes from different classes are entangled with each other, while the supervised manifold learning outcomes are more separable according to the class. Furthermore, although the intrinsic dimension is one, at least five PCs are required to explain 95% variances of the data, indicating inefficiency of the FPCA.
@@ 2.2 Main Method
We see from the examples in Section 2.1 that classification on manifold data can be sufficiently complicated especially under the functional data setting, even if the intrinsic dimension is low and the intrinsic class boundary is simple. To tackle the challenge, our aim is to learn low-dimensional representations of $\{X_{i}\}_{i=1}^{n}$, also termed as manifold learning outcomes, that
@@
It is expected that a better classifier can be constructed based on these low-dimensional representations. The procedure with target (1) alone is termed unsupervised manifold learning, while that with both (1) and (2) is known as supervised manifold learning. Note that for the purpose of classification, the target (2) alone is not sufficient. That is, the classification model trained on manifold learning outcomes would not work well, if the individuals from different classes are separated as much as possible but the geometric structure is fully destroyed. This is because, to predict the label for a future observation, we need to construct a smooth interpolator from the functional space $L^{2}(\mathcal{T})$ to the manifold learning outcome space $\mathbb{R}^{d}$, which performs well only if the target (1) is achieved properly. To illustrate this argument, consider two adjacent points $X_{1}$ and $X_{2}$ with $Y_{1}\neq Y_{2}$. If the corresponding manifold learning outcomes $\widetilde{Z}_{1}$ and $\widetilde{Z}_{2}$ are far away, then it is impossible to accurately predict the manifold learning outcome of $X_{0}$ that is close to $X_{1}$ and $X_{2}$; see Section 4 for theoretical justification. Moreover, ignoring the target (1) is negative for meaningful data visualization. Therefore, we intend to achieve a proper balance between these two targets: The desired manifold learning outcomes gather or disperse according to their classes, in the meantime, the particular manifold structure is maintained as well as possible.
@@
A basic idea in classification is that close or similar individuals should be assigned into the same class. A suitable measure to describe the proximity of two points in a manifold is the geodesic distance $d_{\mathcal{M}}$. Therefore, the aspect of the manifold structure that we intend to preserve is the geodesic distance, which is the same as the idea of the Isomap (Tenenbaum et al.,, 2000). Given pairwise geodesic distances of $\{X_{i}\}_{i=1}^{n}$, low-dimensional representations $\{\widetilde{Z}_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ are obtained by applying the multidimensional scaling (MDS), where the Euclidean distances among the $\widetilde{Z}_{i}$’s are approximately the corresponding geodesic distances among the $X_{i}$’s. Roughly speaking, $\{\widetilde{Z}_{i}\}_{i=1}^{n}$ represents a discrete version of the unwrapped $\mathcal{M}$ embedded into $\mathbb{R}^{d}$.
@@
However, due to the dimension reduction from $\{X_{i}\}_{i=1}^{n}$ to $\{\widetilde{Z}_{i}\}_{i=1}^{n}$ and estimation errors from a finite sample, the (estimated) distribution supports of different classes of $\{\widetilde{Z}_{i}\}_{i=1}^{n}$ in $\mathbb{R}^{d}$ may be nonlinearly entangled, even if the underlying true distributions of different classes of $X$ in $\mathcal{M}$ are easily separated; see Example 2. In this case, a classifier trained on $\{\widetilde{Z}_{i}\}_{i=1}^{n}$ does not accurately approximate the optimal classifier on $\mathcal{M}$. Therefore, we propose penalizing the geodesic distances of individuals of different classes. Specifically, letting $\mathbb{D}_{\xi}$ be the $n\times n$ proximity matrix, we define its $(i,j)$th-entry as
$\mathbb{D}_{\xi}(i,j)=d_{\mathcal{M}}(X_{i},X_{j})+\mathds{1}\{Y_{i}\neq Y_{j}\}\dfrac{\xi}{d_{\mathcal{M}}(X_{i},X_{j})+\sqrt{\xi}}\,,$ (2.1)
@@
where $\mathds{1}\{\cdot\}$ is the indicator function and $\xi>0$ is a tuning parameter. That is, for two individuals from different classes, their geodesic distance is penalized by adding $\xi/\{d_{\mathcal{M}}(X_{i},X_{j})+\sqrt{\xi}\}$. The motivation for using $\xi/\{d_{\mathcal{M}}(X_{i},X_{j})+\sqrt{\xi}\}$ is twofold. First, the penalty is negative correlated with $d_{\mathcal{M}}(X_{i},X_{j})$. This is desirable, because for two points with large $d_{\mathcal{M}}(X_{i},X_{j})$, their classes being different can be largely quantified by their geodesic distance, and there is no need to penalize further on their proximity. Second, the proximity rank for individuals from different classes remains unchanged, i.e., $\mathbb{D}_{\xi}(i,j)<\mathbb{D}_{\xi}(i,k)$ if $d_{\mathcal{M}}(X_{i},X_{j})<d_{\mathcal{M}}(X_{i},X_{k})$, for $Y_{i}\neq Y_{j}$ and $Y_{i}\neq Y_{k}$. These two desired features of $\xi/\{d_{\mathcal{M}}(X_{i},X_{j})+\sqrt{\xi}\}$ properly meet the aforementioned target (1), i.e., alleviating rupture of the original manifold structure (in particular, pairwise geodesic distances). To ensure the target (2), we may simply set a larger $\xi$. The penalized geodesic distance defined in (2.1) distinguishes our method from the supervised Isomap methods in the literature (Vlachos et al.,, 2002; Geng et al.,, 2005; Zhang et al.,, 2018) and may be of interest in its own right.
By applying the MDS on $\mathbb{D}_{\xi}$, we obtain the supervised manifold learning outcomes $\{Z_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$, which are regarded as low-dimensional representations of the functional data that include the label information. In addition, they can be used for visualization if we set $d=2$ or $3$. We assume that $\mathcal{M}$ is characterized by a single coordinate map $\mu:\mathcal{M}\to\mathcal{U}$, where $\mu$ is a diffeomorphism (i.e., a one-to-one differentiable map whose inverse is also differentiable) and $\mathcal{U}$ is an open set in $\mathbb{R}^{d}$. The assumption that $\mathcal{M}$ is covered by a single chart can be restrictive, but is routinely assumed in the literature of manifold learning for tractable theoretical investigation (Tenenbaum et al.,, 2000; Zhang and Zha,, 2004; Chen and Müller,, 2012). The outcomes $\{Z_{i}\}_{i=1}^{n}$ provide an estimator of $\mu$ at $\{X_{i}\}_{i=1}^{n}$. To construct a functional classifier for future observations, we need to develop a global estimator $\widetilde{\mu}$ of $\mu$, which can be achieved by regressing $\{Z_{i}\}_{i=1}^{n}$ on $\{X_{i}\}_{i=1}^{n}$ using a nonparametric smoothing technique. The procedure is also known as the out-of-sample interpolation in the literature on manifold learning. Finally, our proposed functional classifier is defined as $H(\cdot)=H_{d}\{\widetilde{\mu}(\cdot)\}$, where $H_{d}:\mathbb{R}^{d}\to\mathcal{Y}$ is any suitable classifier trained on $\{Z_{i},Y_{i}\}_{i=1}^{n}$. Specifically, we prefer simple geometry-driven classifiers for constructing $H_{d}$, such as the $k$-NN classifier and the SVM, because we expect that the optimal boundaries for separating classes in the manifold can be well unfolded onto $\mathbb{R}^{d}$.

### v2

@@ Abstract
Classification is a core topic in functional data analysis. A large number of functional classifiers have been proposed in the literature, most of which are based on functional principal component analysis or functional regression. In contrast, we investigate this topic from the perspective of manifold learning. It is assumed that functional data lie on an unknown low-dimensional manifold, and we expect that superior classifiers can be developed based on the manifold structure. To this end, we propose a novel proximity measure that takes the label information into account to learn the low-dimensional representations, also known as the supervised manifold learning outcomes. When the outcomes are coupled with multivariate classifiers, the procedure induces a new family of functional classifiers. In theory, we prove that our functional classifier induced by the $k$-NN classifier is asymptotically optimal. In practice, we show that our method, coupled with several classical multivariate classifiers, achieves highly competitive classification performance compared to existing functional classifiers across both synthetic and real data examples. Supplementary materials are available online.
@@ 1 Introduction
Functional data analysis has become increasingly prominent across diverse scientific fields, including medicine, neuroscience, astronomy, chemometrics, etc. As a result, there is a growing demand for sophisticated models and algorithms designed for functional data analysis. A core topic in functional data analysis is classification, i.e., training a functional classifier based on observed data to predict the label for a future functional observation. There exists a fairly rich literature on this topic. Due to the infinite-dimensional nature of functional data, a large number of existing functional classifiers are built on the functional principal component analysis (FPCA), whose theoretical basis is due to the Karhunen–Loève expansion: A second order random process $X$ can be decomposed as
@@
where $\mu(t)=E\{X(t)\}$, $\zeta_{j}$ denote the principal component (PC) scores, $\phi_{j}$ represents the eigenfunctions of the covariance operator, also referred to as the PC basis. A finite number of the PC scores serves as a finite-dimensional approximation of the random process. Delaigle and Hall, (2013) and Galeano et al., (2015) adapted the linear and quadratic discriminant analysis to the PC scores to develop functional linear and quadratic discriminant classifiers. Delaigle and Hall, (2012) proposed to project functional data onto a one-dimensional space spanned by a linear combination of the PC basis, and showed that asymptotic perfect classification can be achieved under specific conditions. Wu and Liu, (2013) treated a finite number of PC scores as predictors and performed the classical support vector machines on the PC scores for functional classification. Dai et al., (2017) proposed to estimate the densities of the PC scores nonparametrically and to utilize the Bayes classifier. Wang et al., 2024b constructed a deep neural network on the PC scores as a functional classifier, and Wang et al., (2023) generalized it to the setting of multidimensional functional data. In addition to FPCA-based methods, other functional classifiers include regression-based methods (Ferraty and Vieu,, 2003; Leng and Müller,, 2006; Preda et al.,, 2007; Araki et al.,, 2009), distance-based methods (Alonso et al.,, 2012; Kraus and Stefanucci,, 2019), and nearest neighbors (Biau et al.,, 2005, 2010); see Wang et al., 2024a for a recent review.
@@
In this article, we propose a novel manifold learning procedure for functional data by taking the label information into account, which induces a new family of functional classifiers. Instead of the FPCA or regression, we start from the assumption that functional data lie on an unknown low-dimensional manifold. This assumption for high-dimensional data forms the basis for manifold learning (Meilă and Zhang,, 2024), and has also been invoked under the functional data setting (Chen and Müller,, 2012; Lin and Yao,, 2021; Tan et al.,, 2024). A motivating example of the manifold assumption is that functional data often exhibit phase variation, which inflates the variance of data and distorts the PCs (Marron et al.,, 2015). To investigate the non-linearity of phase variation, manifold techniques have been developed (Kneip and Gasser,, 1992; Srivastava et al.,, 2011). In addition, Marron and Dryden, (2021) is a relatively new monograph that collects many useful techniques to deal with non-Euclidean data, including manifold data analysis, curve registration, and distance-based classification that are relevant to our work.
@@
Based on the manifold assumption, we propose a novel proximity measure that incorporates the label information to quantify distance between functional data. Not only does the proximity measure penalize for different classes, but it also tries to maintain the underlying manifold structure in terms of the geodesic distance. The latter property is crucial to classification, because the manifold learning step being continuous is necessary for an interpolator to accurately predict the low-dimensional representation of a future observation; see Section 2.2 for more discussion. Once the proximity measure is obtained for all pair-wise individuals, we perform the multidimensional scaling (MDS, Cox and Cox,, 2008) to obtain the low-dimensional manifold learning outcomes. Such outcomes can also be used for data visualization if the target dimension is two or three. Finally, a smooth interpolator accompanied with any reasonable multivariate classifier yields a functional classifier. In theory, we provide the with-in-sample and out-of-sample convergence rates of the estimated coordinate map. As for classification, we also show that our functional classifier coupled with the vanilla $k$-NN classifier is asymptotically optimal. Note that the intrinsic dimension being low is opposite to the number of PCs being high when the manifold is flat. This latter point is a core assumption to prove nice theoretical properties, say, the asymptotic perfect classification, of several functional classifiers (Delaigle and Hall,, 2012; Dai et al.,, 2017; Berrendero et al.,, 2018). Therefore, our investigation of the low intrinsic dimension serves as an interesting complement to the existing literature.
@@
Manifold learning has been an active research field since the seminal works by Tenenbaum et al., (2000) and Roweis and Saul, (2000), which has gained tremendous success in bioinformatics (Moon et al.,, 2018), image processing (Pless and Souvenir,, 2009), etc. In functional data analysis, Chen and Müller, (2012) extended the Isomap (Tenenbaum et al.,, 2000) to functional data and developed the functional manifold component; Tan et al., (2024) extended the parallel transport unfolding (Budninskiy et al.,, 2019) to functional data and conducted a cluster analysis. However, their works belong to unsupervised manifold learning, i.e., no label variable is involved, and thus their methods are not suitable for classification. In contrast, our manifold learning procedure is designed to take the label information into account, also known as supervised manifold learning. There is a vast literature on supervised manifold learning for high-dimensional data; see Chao et al., (2019) for a survey. Also, there exist classifiers based on distances penalized by different classes; see e.g., Ch 11.4 of Marron and Dryden, (2021). Our method may be categorized into Isomap-based methods, but it differs from existing supervised Isomap methods (Vlachos et al.,, 2002; Geng et al.,, 2005; Zhang et al.,, 2018) in two main aspects: Our proximity measure is new, and we are dealing with functional data instead of high-dimensional data.
@@
The rest of this article is organized as follows. In Section 2.1, we first introduce the model and present two examples to show the difficulty of the problem and the motivation of our method. We then introduce our main method in Section 2.2. In Section 3, we discuss a detailed estimation procedure to implement the proposed method, including the selection of tuning parameters. We provide theoretical properties of our estimators in Section 4, and conduct simulation studies and real data applications in Sections 5 and 6, respectively. We conclude in Section 7.
@@ 2.1 Model and Synthetic Examples
It is reasonable to assume that the functional variables from different classes concentrate in different regions of the manifold, i.e., the modes of the distributions $X|Y=y$ are easily separated in $\mathcal{M}$. However, the optimal boundaries for separating classes may be highly nonlinear in the original space, which means that standard linear classification methods may fail. Yet, the optimal boundaries may become simpler if the data are properly embedded in a low-dimensional space. Example 1 illustrates this phenomenon in $\mathbb{R}^{3}$, and Example 2 is a more concrete example showing the classification difficulty under the functional manifold setting.
@@
Figure 2: The panels from left to right show a simulated sample of curves as in Example 2, the first two principal component scores (PC1 and PC2) using the FPCA, the unsupervised manifold learning outcomes $\widetilde{Z}=(\widetilde{Z}_{1},\widetilde{Z}_{2})$ using the FIsomap and the supervised manifold learning outcomes $Z=(Z_{1},Z_{2})$ using our proposed method FSML, respectively. The class $Y=1$ is denoted by solid line (-) and star (*), while $Y=2$ is denoted by dashed line (--) and circle ($\circ$).
@@
Consider a functional time warping model $X(t)=g\{\gamma(t)\}$, for $t\in[0,1]$, which is commonly used for modeling phase variation. We set $\gamma(t)=\int_{0}^{t}\exp\{\theta(s)\}\,ds/\int_{0}^{1}\exp\{\theta(s)\}\,ds$, where $\theta(t)=\theta_{\omega}(t)=\sum_{k=1}^{k_{0}}\sin(k\omega)\sin(\pi kt)/\{(k-8)^{2}+2\}$ with $\omega\sim\textrm{Unif}[0,2\pi]$, for $t\in[0,1]$. The functional variable $X$ can be seen as lying on a manifold of dimension one, as it is generated by a one-dimensional variable $\omega$. Yet, the valued space of $X$ is sufficiently complicated: We can show that $\Theta=\{\theta(\cdot)\mid\omega\in[0,2\pi],k_{0}=\infty\}$, as a subspace of $L^{2}([0,1])$, cannot be covered by any finite-dimensional linear subspace of $L^{2}([0,1])$; see Section A of the Appendix for details.
For the label variable $Y$, we set $Y=1$ if $w\in[0,\pi]$ and $Y=2$ otherwise. The first panel of Figure 2 shows a simulated sample of 200 curves (half in each class) with $g=\sin(2\pi t)$ and $k_{0}=50$. We apply the FPCA to the sample, whose first two principal component scores are shown in the second panel of Figure 2. The manifold learning outcomes using the FIsomap (Chen and Müller,, 2012) and our proposed functional supervised manifold learning (FSML) are shown in the third and fourth panels of Figure 2, respectively. We see that the PC scores and the FIsomap outcomes from different classes are entangled with each other, while the supervised manifold learning outcomes are more separable by class. Furthermore, although the intrinsic dimension is one, at least five PCs are required to explain 95% variances of the data, indicating inefficiency of the FPCA.
Remark 2.
As suggested by a referee, to tackle with functional data with phase variation, one can first perform a curve registration (e.g., the Fisher-Rao registration, Srivastava et al.,, 2011) and then classify the resulting warping functions using standard functional classifiers. In our numerical examples, we show that this strategy does not evidently improve classification accuracy for functional data with phase variation and often worsens classification accuracy for those without phase variation; see Section E of the Appendix for details.
@@ 2.2 Main Method
We see from the examples in Section 2.1 that classification on manifold data can be sufficiently complicated especially under the functional data setting, even if the intrinsic dimension is low and the intrinsic class boundary is simple. To tackle the challenge, our aim is to learn low-dimensional representations of $\{X_{i}\}_{i=1}^{n}$, also termed manifold learning outcomes, that
@@
It is expected that a better classifier can be constructed based on these low-dimensional representations. The procedure with target (1) alone is termed unsupervised manifold learning, while that with both (1) and (2) is known as supervised manifold learning. It is important to note that target (2) alone is insufficient for classification purposes. That is, the classification model trained on manifold learning outcomes would not work well, if the individuals from different classes are separated as much as possible but the geometric structure is significantly compromised. This is because, to predict the label for a future observation, we need to construct a smooth interpolator from the functional space $L^{2}(\mathcal{T})$ to the manifold learning outcome space $\mathbb{R}^{d}$, which performs well only if the target (1) is properly achieved. To illustrate this argument, consider two adjacent points $X_{1}$ and $X_{2}$ with $Y_{1}\neq Y_{2}$. If the corresponding manifold learning outcomes $\widetilde{Z}_{1}$ and $\widetilde{Z}_{2}$ are far away, then it is impossible to accurately predict the manifold learning outcome of $X_{0}$ that is close to $X_{1}$ and $X_{2}$; see Section 4 for theoretical justification. Moreover, two- or three-dimensional manifold learning outcomes that ignore the target (1) render data visualization less useful, because they do not reveal the manifold structure of the covariate. Therefore, we intend to achieve a proper balance between these two targets: The desired manifold learning outcomes gather or disperse according to their classes; meanwhile, the particular manifold structure is maintained as well as possible.
@@
A basic idea in classification is that close or similar individuals should be assigned into the same class. A suitable metric to describe the proximity of two points in a manifold is the geodesic distance $d_{\mathcal{M}}$. Therefore, the aspect of the manifold structure that we intend to preserve is the geodesic distance, which is the same as the idea of the Isomap (Tenenbaum et al.,, 2000). Given pairwise geodesic distances of $\{X_{i}\}_{i=1}^{n}$, low-dimensional representations $\{\widetilde{Z}_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ are obtained by applying the multidimensional scaling (MDS), where the Euclidean distances among the $\widetilde{Z}_{i}$’s approximate the corresponding geodesic distances among the $X_{i}$’s. Roughly speaking, $\{\widetilde{Z}_{i}\}_{i=1}^{n}$ represents a discrete representation of the unwrapped $\mathcal{M}$ embedded into $\mathbb{R}^{d}$.
@@
However, due to the dimension reduction from $\{X_{i}\}_{i=1}^{n}$ to $\{\widetilde{Z}_{i}\}_{i=1}^{n}$ and estimation errors from a finite sample, the (estimated) distribution supports of different classes of $\{\widetilde{Z}_{i}\}_{i=1}^{n}$ in $\mathbb{R}^{d}$ may be nonlinearly entangled, even if the underlying true distributions of different classes of $X$ in $\mathcal{M}$ are easily distinguishable; see Example 2. In such scenarios, a classifier trained on $\{\widetilde{Z}_{i}\}_{i=1}^{n}$ fails to accurately approximate the optimal classifier on $\mathcal{M}$. Therefore, we propose penalizing the geodesic distances of individuals of different classes. Specifically, letting $\mathbb{D}_{\xi}$ be the $n\times n$ proximity matrix, we define its $(i,j)$th-entry as
$\mathbb{D}_{\xi}(i,j)=d_{\mathcal{M}}(X_{i},X_{j})+\dfrac{\xi\cdot\mathds{1}\{Y_{i}\neq Y_{j}\}}{d_{\mathcal{M}}(X_{i},X_{j})+\sqrt{\xi}}\,$, (2.1)
@@
where $\mathds{1}\{\cdot\}$ is the indicator function, and $\xi>0$ is a tuning parameter. That is, for two individuals from different classes, their geodesic distance is penalized by adding $\xi/\{d_{\mathcal{M}}(X_{i},X_{j})+\sqrt{\xi}\}$. The motivation for using $\xi/\{d_{\mathcal{M}}(X_{i},X_{j})+\sqrt{\xi}\}$ is twofold. First, the penalty is negatively correlated with $d_{\mathcal{M}}(X_{i},X_{j})$. This is desirable, because for two points with large $d_{\mathcal{M}}(X_{i},X_{j})$, their classes being different can be largely quantified by their geodesic distance, and there is no need to penalize further on their proximity. Second, the proximity rank for individuals from different classes remains unchanged, i.e., $\mathbb{D}_{\xi}(i,j)<\mathbb{D}_{\xi}(i,k)$ if $d_{\mathcal{M}}(X_{i},X_{j})<d_{\mathcal{M}}(X_{i},X_{k})$, for $Y_{i}\neq Y_{j}$ and $Y_{i}\neq Y_{k}$. These two desired features of $\xi/\{d_{\mathcal{M}}(X_{i},X_{j})+\sqrt{\xi}\}$ properly meet the aforementioned target (1), i.e., alleviating rupture of the original manifold structure (in particular, pairwise geodesic distances). To ensure the target (2), we may simply set a larger $\xi$. The penalized geodesic distance defined in (2.1) distinguishes our approach from the supervised Isomap methods in the literature (Vlachos et al.,, 2002; Geng et al.,, 2005; Zhang et al.,, 2018) and may be of interest in its own right.
[…]